### v1

# ATOM: A Pretrained Neural Operator for Multitask Molecular Dynamics

Abstract
Molecular dynamics (MD) simulations underpin modern computational drug discovery, materials science, and biochemistry. Recent machine learning models provide high-fidelity MD predictions without the need to repeatedly solve quantum mechanical forces, enabling significant speedups over conventional pipelines. Yet many such methods typically enforce strict equivariance and rely on sequential rollouts, thus limiting their flexibility and simulation efficiency. They are also commonly single-task, trained on individual molecules and fixed timeframes, which restricts generalization to unseen compounds and extended timesteps. To address these issues, we propose Atomistic Transformer Operator for Molecules (ATOM), a pretrained transformer neural operator for multitask molecular dynamics. ATOM adopts a quasi-equivariant design that requires no explicit molecular graph and employs a temporal attention mechanism, allowing for the accurate parallel decoding of multiple future states. To support operator pretraining across chemicals and timescales, we curate TG80, a large, diverse, and numerically stable MD dataset with over 2.5 million femtoseconds of trajectories across 80 compounds. ATOM achieves state-of-the-art performance on established single-task benchmarks, such as MD17, RMD17 and MD22. After multitask pretraining on TG80, ATOM shows exceptional zero-shot generalization to unseen molecules across varying time horizons. We believe ATOM represents a significant step toward accurate, efficient, and transferable molecular dynamics models.
1 Introduction
Molecular dynamics (MD) serves as a computational microscope of atomic motion and is now integral to drug discovery and materials science pipelines (Dror et al., 2012; De Vivo et al., 2016). In ab initio molecular dynamics, quantum-mechanical density functional theory (DFT) is used to compute atomic forces, and the resulting equations of motion are integrated to generate high-fidelity trajectories. However, DFT’s computational complexity scales at least cubically with the number of atoms, and relies on double-precision arithmetic that limits GPU acceleration (Kresse & Furthmüller, 1996; Stein et al., 2020; Li et al., 2024).
Neural approaches have recently emerged as a promising solution to the scalability bottleneck. Equivariant architectures, in particular, encode physical symmetries to model interatomic dynamics, achieving ab initio-level accuracy at significantly reduced computational cost (Batzner et al., 2022; Musaelian et al., 2022; Batatia et al., 2022, 2023; Xu et al., 2024). While equivariance is often deemed essential for improving generalization, strict symmetry preservation involves substantial tradeoffs (Xu et al., 2024; Schreiner et al., 2023). Architectures that enforce exact equivariance at every layer often increase computational overhead, restrict model expressivity, and complicate optimization (Fuchs et al., 2020; Brehmer et al., 2023; Elhag et al., 2025). It is unclear whether symmetry constraints can be relaxed without sacrificing accuracy for molecular dynamics.
Furthermore, most existing methods for molecular dynamics are autoregressive, predicting the next state based on the current one (Köhler et al., 2019; Fuchs et al., 2020; Thiemann et al., 2025). Autoregressive approaches often struggle to capture long-horizon temporal dependencies and accumulate error as the prediction horizon grows (Bengio et al., 2015; Bergsma et al., 2023; Taieb & Atiya, 2016). Inference speeds are also constrained by the need for sequential integration, failing to exploit modern, highly parallel compute architectures. One exception is Equivariant Graph Neural Operator (EGNO) (Xu et al., 2024), which models the entire trajectory with neural operator learning. Nevertheless, EGNO enforces strict equivariance and is single-task in nature, i.e., it is trained and evaluated on trajectories of each molecule separately with a fixed time horizon, which limits zero-shot generalization to unseen molecules or timeframes.
Our Main Contributions. In this work, we address the above issues regarding equivariance, autoregression, and zero-shot generalization within a unified framework, which we call Atomistic Transformer Operator for Molecules (ATOM). To this end, we propose a pre-trained neural operator with a transformer backbone for molecular dynamics and introduce a new MD dataset, TG80, which is both chemically diverse and numerically stable for multitask pretraining and benchmarking.
  • Design innovations. ATOM is quasi-equivariant by employing an equivariant lifting layer that produces symmetry-aware features, while allowing subsequent transformer blocks to be unconstrained for flexibility and expressiveness. Unlike autoregressive models, ATOM allows parallel decoding of molecule states across multiple timesteps, directly learning the trajectory operator. By encoding time lags via a novel temporal rotary position embedding, ATOM enhances temporal interpolation and extrapolation, enabling robust predictions across multiple time horizons. Finally, ATOM requires no predefined molecular graph and operates directly on point clouds, naturally accommodating long-range spatial interactions without the need for hand-crafted connectivity.
  • Performance highlights. ATOM sets new state-of-the-art on single-task MD benchmarks. For larger, sparsely connected molecules in MD22, ATOM significantly outperforms existing graph-based baselines by capturing the long-range atomic interactions. In the multitask regime, we pretrain ATOM on TG80 trajectories from multiple molecules and varying timeframes, demonstrating significant zero-shot transfer to both unseen molecules and timesteps, improving existing baselines by 39.75% on average. This achieves performance on par with existing specialized baselines tailored for such molecules and timeframes. To the best of our knowledge, this is the first method that demonstrates such generalization capability in molecular dynamics.
We believe our work represents a shift in molecular dynamics modeling, where we demonstrate the potential of quasi-equivariance designs and zero-shot generalization to out-of-domain systems, which is enabled by the comprehensive TG80 MD dataset.
2 Related work
Equivariant Neural Networks. Equivariance (to transformations such as rotation, reflection, and translation) has emerged as an essential physics-informed prior for deep learning models on molecular data (Bronstein et al., 2021; Duval et al., 2023). Early works employed convolutional approaches to achieve translation equivariance in E(3) (Weiler et al., 2018; Wu et al., 2020) or tensor product attention and spherical harmonics to enforce roto-translational equivariance in SE(3) (Fuchs et al., 2020; Thomas et al., 2018). In contrast, message passing neural network (MPNN) frameworks, such as Equivariant Graph Neural Network (EGNN) and others (Garcia Satorras et al., 2021; Gasteiger et al., 2021; Huang et al., 2022), achieve equivariance by operating on strictly equivariant features, such as inter-node distances and directions. While effective, MPNNs typically assume a fixed molecular graph. This is problematic when the underlying structure contains non-local interactions and dynamic bonding effects (e.g., resonances, transient interactions), which render predefined graphs inaccurate over time (Knutson et al., 2022; Luo et al., 2021). To address this issue, we model molecules as point clouds, with our attention represented as a fully connected graph that allows unrestricted information propagation across the molecule.
Neural Operators. Neural operators are deep learning methods for learning operators between function spaces (Kovachki et al., 2021). A wide variety of architectures have been proposed for such operator learning. Notably, Fourier Neural Operator (FNO) (Li et al., 2021) learns an operator in the Fourier domain, while its derivatives G-FNO (Helwig et al., 2023) and PINO (Li et al., 2023b), respectively, add group equivariance and physics-informed properties. Xu et al. (2024) bridges this framework with molecular dynamics by recasting the task as learning a propagation operator that evolves historical atomic positions into their future configurations. Specifically, EGNO (Xu et al., 2024) is proposed by integrating EGNN and FNO layers to learn dynamic trajectories, capturing both spatial and temporal correlations. Recently, transformer neural operators (Bryutkin et al., 2024; Hao et al., 2023; Li et al., 2023a) have surpassed the performance of FNO in most partial differential equation (PDE) tasks. Notably, OFormer (Li et al., 2023a) uses a linear Galerkin-type attention mechanism, which omits the softmax and instead interprets the latent column vectors as basis functions. General Neural Operator Transformer (GNOT) (Hao et al., 2023) employs a novel subquadratic cross-attention methodology to integrate multiple feature types (e.g., shape and point relationships) into their transformer blocks. With ATOM, we unify the MD problem formulation and temporal discretization approach introduced by EGNO with the increased representational power of transformers in operator settings.
MD Benchmarks. Research on graph machine learning for molecular dynamics suffers from poor benchmarking (Bechler-Speicher et al., 2025). For example, despite the fact that MD17 Benzene exhibits non-physical noise approximately 1000 times higher compared to other compounds (Christensen & von Lilienfeld, 2020), it is still regularly employed to benchmark new models (Bihani et al., 2023; Huang et al., 2022; Liao & Smidt, 2023; Xu et al., 2024). The practical relevance of single-task learning on these datasets is also dubious, as predicting trajectories for molecules with existing numerical solutions offers minimal benefit. We believe the strengths of neural approaches emerge in transfer learning, where models generalize to unseen compounds, thereby circumventing the computational costs associated with explicit numerical simulations. This motivates our development of TG80 to facilitate multitask dynamics learning across molecular systems.
3 Atomistic Transformer Operator for Molecules (ATOM)
In this section, we first introduce the problem formulation (Section 3.1) and then propose the framework of ATOM by introducing the key model and training designs (Section 3.2). We then discuss the multitask pretraining for ATOM and introduce TG80 MD dataset (Section 3.3).
3.1 Problem Formulation
We follow (Xu et al., 2024) to cast molecular dynamics prediction as operator learning. We model a molecule of $N$ atoms as a point cloud in $\mathbb{R}^{3}$, which we denote as $\mathcal{G}^{(t)}$ for a given system state time $t$. In particular, we write $\mathcal{G}^{(t)}=(\mathbf{x}_{i}^{(t)},\mathbf{v}_{i}^{(t)})_{i=1}^{N}$ that represent molecules in terms of the atom positions $\mathbf{x}$ and velocities $\mathbf{v}$. Our objective is to predict a future trajectory $\mathcal{G}^{(t+\Delta t)}$, where $\Delta t\in[0,\Delta T]$.
Similar to (Xu et al., 2024), we focus on predicting the position states only. Let $\mathcal{U}\colon[0,\Delta T]\rightarrow\mathbb{R}^{N\times 3}$ be the trajectory function mapping $\Delta t$ to $U(\Delta t)\in\mathbb{R}^{N\times 3}$ representing molecule positions $\Delta t$ in the future. We assume a solution operator $F^{\dagger}\colon\mathcal{G}^{(t)}\rightarrow\mathcal{U}$ exists which provides the underlying future trajectory given system states at $t$. Thus, the goal of molecular dynamics prediction becomes training a neural operator $F_{\theta}(\mathcal{G}^{(t)})$ to approximate the target trajectory function $F^{\dagger}(\mathcal{G}^{(t)})$: $\min_{\theta}\mathbb{E}_{\mathcal{G}^{(t)}}\mathcal{L}\big(F_{\theta}(\mathcal{G}^{(t)})(t),F^{\dagger}(\mathcal{G}^{(t)})(t)\big)$, for some loss function $\mathcal{L}\colon\mathcal{U}\times\mathcal{U}\rightarrow\mathbb{R}$. Here, expectation is with respect to the different initial states. By discretizing over the temporal domain and considering $L_{2}$ loss, we optimize the neural operator with a discretized temporal sampling of the states:
$\min_{\theta}\frac{1}{P}\sum_{p=1}^{P}\mathbb{E}_{\begin{subarray}{c}\mathcal{G}^{(t)}\\ \end{subarray}}\left\|F_{\theta}\left(\mathcal{G}^{(t)}\right)(\Delta t_{p})-F^{\dagger}\left(\mathcal{G}^{(t)}\right)(\Delta t_{p})\right\|_{2}^{2}.$ (1)
where $\{\Delta t_{1},...,\Delta t_{p}\}$ are discrete timesteps. We replace the true future state $F^{\dagger}(\mathcal{G}^{(t)})(\Delta t_{p})$ with the known future ground truth node positions $\mathbf{x}^{(t+\Delta t_{p})}$ for $\Delta t_{p}\in[0,\Delta T]$.
Single- and multitask. Unlike prior works (Schreiner et al., 2023; Xu et al., 2024), we consider both single-task and multitask settings. Single-task refers to the case where a separate model is independently trained and evaluated on each molecule and fixed timeframes. This corresponds to the conventional practice in molecular dynamics benchmarks. Multitask instead pretrains one unified model on several molecules across varying time lags and evaluates out-of-domain trajectories on unseen molecules, thereby directly testing zero-shot cross-molecule generalization. Under a multitask setting, the objective (1) computes the expectation over trajectories of multiple molecules.
3.2 ATOM Model and Training Design
Here we outline the pipeline of ATOM. At its core is an equivariant lifting layer (Section 3.2.1), which maps atomic positions, velocities and their phase features into a richer embedding space while preserving symmetry under the Euclidean group $E(3)$. The lifted embeddings are then processed by the ATOM attention block, which applies heterogeneous attention over positions, velocities, and phase features with chemical augmentation (Section 3.2.2). To capture temporal dynamics, we incorporate a temporal rotary position embedding (T-RoPE) (Section 3.2.2) that depends only on time lags and is shared across atoms, ensuring translation invariance in time and permutation invariance within each molecule. Finally, to counter numerical noise in training trajectories, we inject randomly sampled position and velocity perturbations during training (Section 3.2.3), which improves robustness and acts as a regularizer against overfitting. The overall pipeline of ATOM is in Figure 1.
Figure 1: ATOM Pipeline. We pretrain ATOM on the TG80 dataset across multiple molecules with stochastic time lags. At inference, ATOM takes a query molecule and timestamps and directly outputs corresponding molecular states.
3.2.1 E(3) Equivariant Lifting
To model atomic states in a symmetry-respecting way, each atom is encoded with its 3D position and velocity, augmented with their norms: $\mathbf{x}=(x,y,z,\sqrt{x^{2}+y^{2}+z^{2}}),\,\mathbf{v}=(v_{x},v_{y},v_{z},\sqrt{v_{x}^{2}+v_{y}^{2}+v_{z}^{2}})$. To construct higher-dimensional features that remain consistent with $E(3)$ symmetry, we apply equivariant lifting that maps the inputs through learnable functions that preserve group actions. Specifically, we use $E(3)$-equivariant linear layers (Geiger & Smidt, 2022) that lifts the position and velocity vectors to a feature space. The resulting features satisfy the equivariance constraints by construction. We further construct phase space of each atom by augmenting the position and velocity vectors with atomic number, which is subsequently processed by a learnable equivariant layer to obtain a lifted representation. The final lifted embedding for a molecule is given by $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ corresponding to position, velocity and phase features. The second dimension aggregates nodes and time for attention and $d_{v}$ is the embedding space dimension.
We highlight that after the equivariant lifting layer, we do not enforce equivariance in the subsequent Transformer blocks. This relaxation leads to improved performance compared to fully equivariant designs and, show robustness to random rotations of the trajectories compared to non-equivariant baselines (see Section 4.4).
3.2.2 ATOM Heterogeneous Temporal Attention
We employ a heterogeneous temporal attention mechanism to enable mixing between multiple features $({\mathbf{X}},{\mathbf{V}},{\mathbf{Z}})\in\mathbb{R}^{3\times NP\times d_{v}}$ across spatial and temporal dimensions. We use the phase space embedding $\mathbf{Z}$} as the query and attend to the key-value pairs formed from all features ${\mathbf{X}},{\mathbf{V}},{\mathbf{Z}}\in\mathbb{R}^{NP\times d_{v}}$. In Figure 7, we show that this improves performance by 6.36% over standard self-attention for single-task prediction. In addition, to encode temporal information, we introduce Temporal RoPE (T-RoPE), an adaptation of RoPE (Su et al., 2023) to irregular time lags by driving the phases with timestamps built from per-step increments $\{\Delta t\}$.
Let the hidden dimension per head be $d_{h}$ (even). We define frequencies $\omega_{k}=b^{-2k/d_{h}}$ for $k=0,\dots,d_{h}/2-1$. Given per-step time increments $\{\Delta t_{p}\}_{p=1}^{P}$, we build timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$, and assign a single rotation to all $N$ atoms at timestep $p$: $\mathbf{R}_{p}=\operatorname{diag}\left(\mathbf{R}(\theta_{p,0}),\dots,\mathbf{R}(\theta_{p,d_{h}/2-1})\right)\in\mathbb{R}^{d_{h}\times d_{h}}$, where $\theta_{p,k}=\frac{\omega_{k}}{\tau}\,\bigl(t_{p}-t_{0}\bigr)$ and $\mathbf{R}(\theta)\in\mathbb{R}^{2\times 2}$ is the rotation matrix with angle $\theta$ and $\tau>0$ is a timescale hyperparameter. Suppose the query molecule state at time $p$ is given as $\mathbf{Q}_{p}\in\mathbb{R}^{N\times d_{h}}$ and key molecule state at time $p^{\prime}$ is $\mathbf{K}_{p^{\prime}}\in\mathbb{R}^{N\times d_{h}}$. We apply $\mathbf{R}_{p},\mathbf{R}_{p^{\prime}}$ to $\mathbf{Q}_{p},\mathbf{K}_{p^{\prime}}$ respectively so that the rotary dot product $\mathbf{Q}_{p}\mathbf{R}_{p}(\mathbf{K}_{p^{\prime}}\mathbf{R}_{p^{\prime}})^{\top}$ depends only on the time interval $t_{p^{\prime}}-t_{p}$. This makes attention translation invariant in time, which allows for interpolation and extrapolation across irregular increments $\{\Delta t_{p}\}$. In addition, sharing the same $\mathbf{R}_{p}$ across all $N$ atoms in a molecule ensures permutation-invariance within a timestep. For aggregated query and key matrices $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{NP\times d_{h}}$, we denote the application of temporal Rotary Position Embedding (RoPE) across $P$ timesteps and $N$ atoms as $\operatorname{T-RoPE}(\mathbf{Q}),\operatorname{T-RoPE}(\mathbf{K})\in\mathbb{R}^{NP\times d_{h}}$.
Specifically, a single-head attention layer of ATOM computes
$\sum_{\mathbf{F}\in\{{\mathbf{X}},{\mathbf{V}},{\mathbf{Z}}\}}\gamma_{\mathbf{F}}\,\operatorname{softmax}\Big(\frac{\operatorname{T-RoPE}(Q({\mathbf{Z}}))\,\operatorname{T-RoPE}(K(\mathbf{F}))^{\top}}{\sqrt{d_{h}}}\Big)V(\mathbf{F}),$
where $Q(\cdot),K(\cdot),V(\cdot)$ represent the query, key and value projections. We introduce learnable weights $\gamma_{\mathbf{F}}$ to modulate the relative importance of each feature. In Section G.1, we show that heterogeneous attention is equivalent to a kernel integral operator and discuss the properties of the kernel.
3.2.3 Training with Label Noise Regularization
Many DFT datasets are inherently noisy (Christensen & von Lilienfeld, 2020), and MD models can overfit to this noise. Motivated by the regularization effect of label noise (Damian et al., 2021; HaoChen et al., 2020), we augment the observed node positions $\mathbf{x}$ and velocities $\mathbf{v}$ by random Gaussian noise $\boldsymbol{\xi}_{\mathbf{x}},\boldsymbol{\xi}_{\mathbf{v}}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ during training. Let ${\mathcal{G}}^{(t)}_{\boldsymbol{\xi}}=(\mathbf{x}_{i}^{(t)}+\boldsymbol{\xi}_{x,i},\mathbf{v}^{(t)}+\boldsymbol{\xi}_{v,i})$ be the noised initial state at time $t$. We minimize the following regularized loss
$\min_{\theta}\frac{1}{P}\sum_{p=1}^{P}\mathbb{E}_{\mathcal{G}^{(t)},\boldsymbol{\xi},\boldsymbol{\xi}_{x}^{p}}\left\|F_{\theta}\left(\mathcal{G}^{(t)}_{\boldsymbol{\xi}}\right)(\Delta t_{p})-(\mathbf{x}^{(t+\Delta t_{p})}+\boldsymbol{\xi}_{x}^{p})\right\|_{2}^{2}.$
A similar strategy has also appeared in graph neural network (GNN)-based MD models and neural operator pretraining (Dauparas et al., 2022; Zhou et al., 2024a; Hao et al., 2024). We only apply noise augmentation during training and evaluate on the unperturbed ground-truth trajectories.
3.3 Multitask ATOM Pretraining and TG80 Dataset
This section adapts ATOM for the multitask setting, where the aim is to predict future trajectories for unseen molecules. In order to more effectively distinguish molecules, we construct a radius graph of 1.6 Å based on atomic positions, and apply random walk positional encoding (Ma et al., 2023; Lobato et al., 2021) to augment the phase vector $\mathbf{z}$. We describe the process in detail in Section D.2 and highlight that such a graph depends only on atomic positions, not chemical bonds.
During multitask training, each mini-batch contains trajectories from multiple molecules. In addition, we perform random sampling for the time lags $\Delta t$ from a log-uniform distribution between $\Delta t_{\rm min}$ and $\Delta T$, namely $\Delta t\sim{\rm LogUnif}(\Delta t_{\rm min},\Delta T)$. This aims to enhance the robustness of interpolation and extrapolation in the temporal domain, a consideration that has been similarly explored in (Schreiner et al., 2023). Let $\mathcal{M}$ denote the set of training molecules and let $\mathcal{G}_{m}^{(t)}$ represent the state of molecule $m\in\mathcal{M}$ at timestamp $t$. We can write the pretraining multitask objective as
$\min_{\theta}\;\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\mathbb{E}_{\mathcal{G}^{(t)}_{m},\,\Delta t\sim\mathrm{LogUnif}(\Delta t_{\min},\,\Delta T)}\left\|F_{\theta}\!\left(\mathcal{G}^{(t)}_{m},\Delta t\right)(\Delta t)-\mathbf{x}^{(t+\Delta t)}_{m}\right\|_{2}^{2},$
where we take expectation with respect to initial states of multiple molecules in the training set, as well as the time lags. Here, we emphasize ATOM also takes a time lag, $\Delta t$, as input.
TG80 Dataset. To facilitate pretraining of our neural operator, we introduce TG80, a superset of the MD17 dataset. The initial seed set comprises 40 molecules: 8 MD17 compounds and 32 additional drug-like molecules selected through expert review. We then augment the seed molecules with structurally similar molecules from the PubChem dataset of 173 million compounds (Bolton et al., 2011).
Figure 2: Construction of TG80 from an initial seed using the PubChem database.
Accepted candidates had an ECFP-4 Tanimoto similarity between 0.875 and 0.925 to at least one seed molecule, and no more than 0.80 similarity to previously accepted molecules, alongside other criteria detailed in Section C.4 (Landrum et al., 2025; Rogers & Hahn, 2010; Rogers & Tanimoto, 1960). These thresholds follow common practice in the literature, balancing diversity while avoiding collapse into overly narrow chemical subspaces (Matter, 1997; Menke et al., 2021; Eastman et al., 2023; Harper et al., 2004; Zhang et al., 2023).
We generate all trajectories using ORCA V6.01 (Neese, 2022) with the PBE functional (Perdew et al., 1996), def2-SVP basis set (Weigend & Ahlrichs, 2005), $\Delta$4 dispersion corrections (Caldeweyher et al., 2019, 2020; Wittmann et al., 2024) at one femtosecond resolution, 300K temperature, in vacuum. This resembles an enhanced RMD17, with more modern dispersion corrections to improve stability and allow for a larger step size (Christensen & von Lilienfeld, 2020). As a result, TG80 exhibits more diverse dynamics and improved numerical stability, with no compound exceeding 50 Å center-of-mass drift in Figure 9111Simulations ran on 32 AMD EPYC 7543 cores with 256 GB RAM per molecule, totalling 806,400 CPU-hours (quoted market cost USD 150 000)..
4 Experiment Results
Metrics. We use State-to-trajectory ($\operatorname{S2T}$) and state-to-state ($\operatorname{S2S}$) error to evaluate ATOM (Xu et al., 2024). Specifically, $\operatorname{S2T}=\frac{1}{P}\sum_{p=1}^{P}\|\hat{\mathbf{x}}_{p}-\mathbf{x}_{p}\|_{2}^{2}$, measures the average discrepancy between the predicted $\hat{\mathbf{x}}$ and ground-truth positions $\mathbf{x}$ across entire trajectories, while $\operatorname{S2S}=\|\hat{\mathbf{x}}_{P}-\mathbf{x}_{P}\|_{2}^{2}$, quantifies the error at the final predicted timestep.
Baselines. For comparison, we include a range of classic to state-of-the-art baselines, including Radial Field (RF) (Köhler et al., 2019), Tensor Field Networks (TFN) (Thomas et al., 2018), SE(3) Transformer (SE(3)-Tr.) (Fuchs et al., 2020), E(n) equivariant graph neural networks (EGNN) (Garcia Satorras et al., 2021), and EGNO (Xu et al., 2024). Our EGNN baselines are EGNN-Rollout (EGNN-R), which predicts timesteps autoregressively, and EGNN-Sequential (EGNN-S), which uses the output of each GNN as the prediction of a given frame. We set all baseline hyperparameters following previous works (Xu et al., 2024, 2022; Shi et al., 2021) and tune ATOM and EGNO hyperparameters as in Table 16 and Table 17.
Training setups. For training of ATOM and EGNO, we consider two temporal discretization strategies in selecting the timestamps $t_{p}=t+\sum_{r=1}^{p}\Delta t_{r}$: (1) Uniform discretization selects $t_{p}=t+p/P\Delta T$ and (2) Tail discretization selects $t_{p}=t+\overline{\Delta}+p/P(\Delta T-\overline{\Delta})$ for a lag $\overline{\Delta}\in[0,\Delta T]$. In the main paper, we present experiment results with uniform discretization and include the results with tail discretization in Appendix E. We perform early stopping on the lowest S2S validation loss checkpoint and report results as mean $\pm 2\sigma$ over three training runs. All experiments are run on an NVIDIA® RTX 5080 with wall-clock time and FLOP utilization detailed in Table 14.
4.1 Single-task Learning
We benchmark on the MD17, RMD17, and MD22 DFT MD trajectory datasets (Chmiela et al., 2017; Christensen & von Lilienfeld, 2020; Chmiela et al., 2023). We partition the trajectories into train/validation/test splits of sizes 500/2000/2000, set $\Delta T=3000$ and $P=8$, and train for 2500 epochs following (Xu et al., 2024). For the performance on MD17 ( Table 1), we directly quote the results from (Xu et al., 2024) except for EGNO. We design ATOM to have six transformer blocks with a hidden size of 256.
MD17 and RMD17. As shown in Table 1, ATOM compares favorably with state-of-the-art (SOTA) baselines on MD17 dataset, yielding average reductions of 14.96% (S2S mean squared error (MSE)) and 8.3% (S2T MSE) on average [^2] We exclude benzene from the table due to the previously discussed high numerical noise.. In Table 9 (Appendix E.2), we benchmark ATOM on RMD17, and observe similarly competitive performance against EGNO.
Figure 3: Docosahexaenoic acid (DHA)
MD22. To evaluate performance on larger molecules, we consider Ac-Ala3-NHMe (20 heavy atoms), docosahexaenoic acid (DHA with 24 heavy atoms), and stachyose (45 heavy atoms) from the MD22 dataset (Chmiela et al., 2023). ATOM remains competitive on these systems; whereas EGNO fails to converge (Table 2). We attribute this discrepancy to differing inductive biases: GNNs such as EGNO restrict message passing to a predefined bond or radius graph and can therefore under-represent long-range, non-bonded steric and electrostatic interactions that dominate the behavior of large, sparsely connected molecules (Alon & Yahav, 2021; Kosmala et al., 2023). This explains the poor performance of EGNO on DHA, a prototypical sparse molecule (Nv et al., 2003) (Figure 3), compared to its relatively stronger results on the densely bonded stachyose. In contrast, the fully connected point-cloud attention in ATOM imposes no prior on interaction ranges, allowing simultaneous learning of both local and global interactions.
Table 1: Single-task MSE ($\times 10^{-2}$) on MD17. Upper part: S2S MSE. Lower part: S2T MSE.
Aspirin Ethanol Malonaldehyde Naphthalene Salicylic Toluene Uracil
RF $10.94{\scriptstyle\pm 0.02}$ $4.64{\scriptstyle\pm 0.02}$ $13.93{\scriptstyle\pm 0.06}$ $0.50{\scriptstyle\pm 0.02}$ $1.23{\scriptstyle\pm 0.04}$ $10.93{\scriptstyle\pm 0.08}$ $0.64{\scriptstyle\pm 0.02}$
TFN $12.37{\scriptstyle\pm 0.36}$ $4.81{\scriptstyle\pm 0.08}$ $13.62{\scriptstyle\pm 0.16}$ $0.49{\scriptstyle\pm 0.02}$ $1.03{\scriptstyle\pm 0.04}$ $10.89{\scriptstyle\pm 0.02}$ $0.84{\scriptstyle\pm 0.04}$
SE(3)-Tr. $11.12{\scriptstyle\pm 0.12}$ $4.74{\scriptstyle\pm 0.02}$ $13.89{\scriptstyle\pm 0.04}$ $0.52{\scriptstyle\pm 0.02}$ $1.13{\scriptstyle\pm 0.04}$ $10.88{\scriptstyle\pm 0.12}$ $0.79{\scriptstyle\pm 0.04}$
EGNN $14.41{\scriptstyle\pm 0.30}$ $4.64{\scriptstyle\pm 0.04}$ $13.64{\scriptstyle\pm 0.02}$ $0.47{\scriptstyle\pm 0.04}$ $02{\scriptstyle\pm 0.04} 11.$$1.78{\scriptstyle\pm 0.14}$ $0.64{\scriptstyle\pm 0.02}$
EGNN-R $9.96{\scriptstyle\pm 0.14}$ $4.61{\scriptstyle\pm 0.01}$ $13.04{\scriptstyle\pm 0.03}$ $0.44{\scriptstyle\pm 0.05}$ $0.96{\scriptstyle\pm 0.01}$ $0.19{\scriptstyle\pm 0.15} 1.1$$11{\scriptstyle\pm 0.04}$
EGNN-S $10.25{\scriptstyle\pm 0.09}$ $4.61{\scriptstyle\pm 0.01}$ $13.06{\scriptstyle\pm 0.01}$ $0.53{\scriptstyle\pm 0.01}$ $1.06{\scriptstyle\pm 0.05}$ $10.83{\scriptstyle\pm 0.09}$ $0.62{\scriptstyle\pm 0.01}$
EGNO $9.64{\scriptstyle\pm 0.15}$ $4.57{\scriptstyle\pm 0.01}$ $\mathbf{12.92}{\scriptstyle\pm 0.00}$ $\mathbf{0.39}{\scriptstyle\pm 0.00}$ $0.89{\scriptstyle\pm 0.01}$ $11.00{\scriptstyle\pm 0.00}$ $\mathbf{0.58}{\scriptstyle\pm 0.02}$
ATOM $\mathbf{6.82}{\scriptstyle\pm 0.06}$ $\mathbf{3.52}{\scriptstyle\pm 0.04}$ $14.72{\scriptstyle\pm 0.01}$ $0.50{\scriptstyle\pm 0.00}$ $\mathbf{0.88}{\scriptstyle\pm 0.01}$ $\mathbf{4.66}{\scriptstyle\pm 0.21}$ $0.63{\scriptstyle\pm 0.00}$
EGNN-R $7.35{\scriptstyle\pm 0.19}$ $3.21{\scriptstyle\pm 0.00}$ $\mathbf{10.75}{\scriptstyle\pm 0.04}$ $\mathbf{0.34}{\scriptstyle\pm 0.06}$ $1.09{\scriptstyle\pm 0.12}$ $4.53{\scriptstyle\pm 0.08}$ $0.89{\scriptstyle\pm 0.02}$
EGNN-S $9.01{\scriptstyle\pm 0.34}$ $3.21{\scriptstyle\pm 0.00}$ $11.20{\scriptstyle\pm 0.03}$ $0.42{\scriptstyle\pm 0.01}$ $1.41{\scriptstyle\pm 0.00}$ $4.86{\scriptstyle\pm 0.04}$ $0.65{\scriptstyle\pm 0.01}$
EGNO $9.64{\scriptstyle\pm 0.15}$ $4.57{\scriptstyle\pm 0.01}$ $12.92{\scriptstyle\pm 0.00}$ $0.39{\scriptstyle\pm 0.00}$ $0.90{\scriptstyle\pm 0.01}$ $10.99{\scriptstyle\pm 0.00}$ $\mathbf{0.58}{\scriptstyle\pm 0.02}$
ATOM $\mathbf{5.62}{\scriptstyle\pm 0.05}$ $\mathbf{2.62}{\scriptstyle\pm 0.04}$ $12.49{\scriptstyle\pm 0.01}$ $0.43{\scriptstyle\pm 0.00}$ $\mathbf{0.86}{\scriptstyle\pm 0.01}$ $\mathbf{2.27}{\scriptstyle\pm 0.10}$ $0.61{\scriptstyle\pm 0.00}$
4.2 Multitask Learning on TG80
We pretrain ATOM on TG80, scaling to six attention blocks with a hidden size of 256. We select stochastic horizons $\Delta T\sim\operatorname{LogUnif}(8,24\,000)$ and use a five-fold, cluster-based cross-validation. Specifically, we compute ECFP-4 fingerprints (Rogers & Hahn, 2010), embed them using UMAP (McInnes et al., 2018), and apply agglomerative clustering (Ward, 1963) to partition compounds into ten disjoint clusters. The folds are then formed by holding out clusters, ensuring that the train/validation/test sets occupy distinct regions of chemical space. This cluster-wise protocol minimizes leakage and more closely reflects the prospective scientific setting in which models must generalize to unseen molecules. Cluster-based approaches present more challenging generalization problems than random splits or common chemical-scaffold-based splits (Guo et al., 2024). In Appendix E.3, we also consider pre-training on a standard random split of molecules.
Table 3 benchmarks ATOM by assessing both in-distribution (ID) and out-of-domain (OOD) S2T performance. For the in-distribution setting, we train, validate, and test on molecules from the same cluster. We observe that ATOM outperforms existing baselines by an average of 83.96% in terms of S2T MSE. We then assess out-of-domain (OOD) generalization performance by predicting the dynamics of unseen compounds drawn from disjoint clusters. Under OOD settings, ATOM nearly halves the S2T MSE of EGNO, with an average improvement of $39.74\%$ across five cluster splits. Notably, OOD ATOM beats ID EGNO performance in four of five folds. This striking zero-shot generalization, realized without any exposure to the test molecules, confirms that ATOM uniquely learns robust, transferable knowledge of molecular dynamics. In Appendix E.4, we show similar outperformance in S2S prediction. In Appendix F.2, we show that the significantly improved multitask performance comes with a modest overhead in training time and in inference latency.
Table 2: Single-task MSE ($\times 10^{-2}$) on MD22. Upper: S2S. Lower: S2T
Ac-Ala3-NHME DHA Stachyose
EGNO $357.89{\scriptstyle\pm 3.94}$ $178.39{\scriptstyle\pm 4.91}$ $42.11{\scriptstyle\pm 0.10}$
ATOM $\mathbf{9.65}{\scriptstyle\pm 0.75}$ $\mathbf{10.60}{\scriptstyle\pm 1.11}$ $\mathbf{21.25}{\scriptstyle\pm 4.20}$
Gap $97.30\% +$$94.06\% +$$49.54\%$
EGNO $232.40{\scriptstyle\pm 6.75}$ $116.45{\scriptstyle\pm 3.34}$ $30.84{\scriptstyle\pm 0.03}$
ATOM $\mathbf{7.55}{\scriptstyle\pm 0.42}$ $\mathbf{9.66}{\scriptstyle\pm 1.16}$ $\mathbf{18.13}{\scriptstyle\pm 3.78}$
Gap $96.75\% +$$91.70\% +$$41.22\%$
Table 3: Multi-task S2T MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
Cluster 1 Cluster 2 Cluster 3 Cluster 4 Cluster 5
ID EGNO $44.23{\scriptstyle\pm 0.68}$ $95.52{\scriptstyle\pm 0.73}$ $141.16{\scriptstyle\pm 0.21}$ $150.92{\scriptstyle\pm 0.11}$ $107.47{\scriptstyle\pm 0.36}$
ATOM $\mathbf{9.71}{\scriptstyle\pm 0.75}$ $\mathbf{18.26}{\scriptstyle\pm 1.58}$ $\mathbf{16.82}{\scriptstyle\pm 1.46}$ $\mathbf{16.93}{\scriptstyle\pm 3.65}$ $\mathbf{17.20}{\scriptstyle\pm 0.46}$
Gap $78.04\%$ $\% 88.0988.09\%$ $88.78\%$ $83.99\%$
OOD EGNO $45.95{\scriptstyle\pm 0.80}$ $115.43{\scriptstyle\pm 13.23}$ $151.74{\scriptstyle\pm 0.57}$ $63.90{\scriptstyle\pm 0.69} 11$$13.68{\scriptstyle\pm 2.50}$
EGNN-S $45.44{\scriptstyle\pm 0.57}$ $7386.15{\scriptstyle\pm 6931.89}$ $152.72{\scriptstyle\pm 0.83}$ $464.22{\scriptstyle\pm 509.48}$ $114.30{\scriptstyle\pm 0.79}$
EGNN-R $44.88{\scriptstyle\pm 0.68}$ $109.62{\scriptstyle\pm 1.92}$ $148.05{\scriptstyle\pm 0.70}$ $61.54{\scriptstyle\pm 0.68} 11$$10.10{\scriptstyle\pm 0.96}$
ATOM $\mathbf{35.05}{\scriptstyle\pm 0.97}$ $\mathbf{106.99}{\scriptstyle\pm 104.64}$ $\mathbf{60.95}{\scriptstyle\pm 4.86}$ $\mathbf{66.68}{\scriptstyle\pm 0.96}$ $\mathbf{47.49}{\scriptstyle\pm 1.59}$
Gap $21.93\%$ $2.40\%$ $58.83\%$ $58.71\%$ $56.88\%$
4.3 Temporal Gap and Timestep Invariance Properties
$\Delta T$ Invariance. We evaluate the performance of ATOM with varying $\Delta T$. We compare ATOM, EGNO, and EGNN on S2T MSE by fixing $P=8$ and sweeping $\Delta T$ logarithmically from 10 to 10 000 on a pretrained in-distribution multitask model (on Cluster 1). In Figure 5, we show that ATOM maintains its extrapolation advantage across the range compared to EGNO, particularly at larger $\Delta T$. Ablating T-RoPE (NoPE) removes such advantage by exihibiting EGNO-like error trend with substantially higher MSE. This underscores T-RoPE’s role in stable time-gap extrapolation.
$P$ invariance. Following the discretization invariance in neural operators, we expect ATOM and EGNO models to show consistent MSE as $P$ varies under uniform discretization (Kovachki et al., 2021). Figure 5 confirms such a conjecture by showing that multitask ATOM pretrained at $P=8$ maintain constant S2T MSE as $P$ ranges from 4 to 24.
Figure 4: Pretrained (ID) multitask S2T MSE across $\Delta T$ values.
Figure 5: ATOM and EGNO are discretization invariant, showing stable S2T MSE.
4.4 Ablation studies
We perform extensive ablations to assess each design choice in ATOM. For single-task performance (Fig. 7) and multitask performance (Fig. 7), we independently toggle components and measure their contributions. Our analysis focuses on equivariant lifting, T-RoPE, label-noise regularization, heterogeneous attention, and random-walk positional encoding (under multitask pretraining).
Table 4: S2T MSE ($\times 10^{-2}$) of a fixed input frame rotated and unrotated by an $SO(3)$ matrix.
ATOM No equivariant Lift
Unrotated $6.76{\scriptstyle\pm 0.69}$ $33.44{\scriptstyle\pm 23.42}$
Rotated $73.04{\scriptstyle\pm 27.01}$ $660.97{\scriptstyle\pm 945.86}$
Gap $66.28{\scriptstyle\pm 26.32}$ $627.53{\scriptstyle\pm 926.53}$
Equivariant lifting. We assess the quasi-equivariant design against fully equivariant and non-equivariant alternatives. As shown in Figure 7, replacing the equivariant lifting introduced in Section 3.2.1 with standard linear layers (no equivariant lifting) markedly degrades the performance of ATOM, increasing S2T MSE by 22.48. Consistently, Section 4.4 shows that non-equivariant lifting is vulnerable under $SO(3)$ rotations of the trajectories, whereas the quasi-equivariant ATOM remains comparatively robust. Notably, the fully-equivariant variant of ATOM, described in appendix Section D.1, also underperforms ATOM in both single-task (Figure 7) and multitask (Figure 7) settings, with the gap exaggerated in the multitask setting. This aligns with recent findings in relaxed equivariance, suggesting that strict equivariance can restrict the model capacity and complicate the optimization process (Elhag et al., 2025).
Heterogeneous attention. We evaluate the contribution from heterogeneous temporal attention by substituting with a standard self-attention on the phase space features. This replacement reduces accuracy, increasing S2S MSE by 0.47.
Temporal Rotary Position Embedding (T-RoPE). In the single-task regime with fixed $\Delta T$ (Figure 7), T-RoPE contributes little to the performance of ATOM, as it effectively reduces to a constant rotational shift. By contrast, with stochastic $\Delta T$, disabling T-RoPE (NoPE) increases MSE by 1.07, consistent with ATOM leveraging the $\tau$ parameter to encode variable time gaps (Figure 7). An EGNO-style sinusoidal positional encoding produces a similar performance degradation.
Label noise regularization. We also test the utility of label noise regularization as in Section 3.2.3. From Figure 7, we observe that removing augmented noise from the position and velocity features increased S2T MSE by 1.21. For the multitask ablation on TG80, we suppress label noise regularization, as the dataset is designed to be numerically stable with small noise.
RWPE. We assess random-walk positional encoding (RWPE) in the multitask pretraining. Figure 7 indicates that RWPE facilitates molecule identification, yielding improved multitask performance.
Figure 6: ATOM ablation on MD17 Aspirin.
Figure 7: ATOM ablation on TG80 Cluster 1.
5 Conclusions
In this work, we demonstrate that carefully designed transformer neural operators enable zero-shot generalization to unseen chemical dynamics. Our experiments on MD17 demonstrate continued good single-task performance, and we present the first molecular neural operator that can successfully learn large molecule dynamics using MD22. Our multitask experiments show that our method learns transferable dynamics knowledge, even without explicit graph representations. In combination with our TG80 dataset, we provide a large-scale open-source benchmark and baselines to evaluate future models and spur further operator research with concrete scientific applicability.
Limitations
[…]
Reproducibility Statement
We provide experiment details, such as choice of hyperparameters and other training configurations in Appendix F. In addition, we will release the TG80 dataset upon acceptance under MIT license for reproducibility.
Appendices
Appendix A LLM Acknowledgment
The authors acknowledge the use of LLM for grammar corrections and for improving the clarity of the manuscript. The LLM was not used for generating original ideas, content, or experimental results. All conceptual contributions, analyses, and conclusions presented in this work are entirely from the authors.
Appendix B Background
This section provides an introduction to the preliminaries of group theory.
B.1 Groups
A group $(G,\circ)$ consists of a non-empty set $G$ and a binary operation $\circ:G\times G\to G$ satisfying the following axioms:
  1. 1.
    Closure: For all $a,b\in G$, the result of the operation $a\circ b$ is also in $G$: $a\circ b\in G$.
  2. 2.
    Identity Element: There exists an element $e\in G$ such that, for all $a\in G$, $a\circ e=e\circ a=a$.
  3. 3.
    Associativity: For all $a,b,c\in G$, $(a\circ b)\circ c=a\circ(b\circ c)$.
  4. 4.
    Inverses: For each $a\in G$, there exists an element $a^{-1}\in G$ such that $a\circ a^{-1}=a^{-1}\circ a=e$.
In general, not all groups are abelian. That is, the binary operation $\circ$ does not necessarily commute: $g\circ h=h\circ g,\forall g,h\in G.$
B.2 Group Representations
A group representation is a homomorphism $\rho:G\to GL(V)$ that assigns an $n\times n$ matrix to each group element $g\in G$, realizing it as a linear transformation. Representations must preserve the binary operation for all members of the group $G$ such that:
$\rho(g\circ h)=\rho(g)\rho(h),\quad\forall g,h\in G.$
A representation $\rho(g)$ is reducible if it can be represented as the direct sum of other representations:
$\rho(g)=\rho_{1}(g)\oplus\rho_{2}(g),\quad\forall g\in G.$
For example, a reducible $4\times 4$ representation of $SU(2)$ can be decomposed into two $2\times 2$ sub-representations:
$\rho(g)=\begin{bmatrix}\rho_{1}(g)&0\\ 0&\rho_{2}(g)\end{bmatrix},\quad\forall g\in SU(2),$
where $\rho_{1}(g)$ and $\rho_{2}(g)$ are the following irreducible representations of $SU(2)$:
$\rho_{1}(g)=\begin{bmatrix}e^{i\theta}&0\\ 0&e^{-i\theta}\end{bmatrix},\quad\rho_{2}(g)=\begin{bmatrix}e^{i\phi}&0\\ 0&e^{-i\phi}\end{bmatrix}$.
By contrast, irreducible representations or irreps cannot be represented as such a direct sum. Formally, they have no non-trivial invariant subspaces $W\subset V$ such that $\rho(g)W\subset W,\forall g\in G$.
Representing inputs as irreps ensures equivariance by constraining each feature to transform predictably under group actions. Given $V=\bigoplus_{i}V_{i}$ with irreps $V_{i}$, the transformation of an input $x\in V$ under $g\in G$ is:
$\rho(g)x=\bigoplus_{i}\rho_{i}(g)x_{i}.$
Each component $x_{i}$ transforms independently according to $\rho_{i}$, preserving symmetry. Scalars remain invariant, while vectors rotate according to standard representations. This decomposition prevents the mixing of differently transforming features, ensuring that all subsequent operations, linear or non-linear, respect the group’s symmetry, thereby maintaining equivariance throughout the network.
Intuitively, the tensor products capture interactions between features in a manner akin to multiplication, producing a higher-dimensional representation. Crucially, this new representation is reducible, so we may decompose it into irreps:
$V\otimes V\cong\bigoplus_{k}V_{k}.$
It is this decomposition that allows the network to project onto individual irreps, achieving non-trivial feature mixing whilst preserving symmetry constraints.
Appendix C Datasets
We present a visualization of a sample trajectory of uracil from three datasets in Figure 8.
Figure 8: 3000 timesteps of uracil trajectory from MD17, RMD17, and TG80.
C.1 Licences
Table 5: Dataset sources and licenses. We release TG80 under the MIT license.
Dataset Source License
MD17 https://www.sgdml.org/ CC BY 4.0
RMD17 https://archive.materialscloud.org/record/2020.82 CC Zero V1.0 Universal
MD22 https://www.sgdml.org/ CC BY 4.0
TG80 To be released at URL MIT
C.2 Model Inputs and the Dataloader
Our compound representations follow (Shi et al., 2021; Xu et al., 2024). We model hydrogen atoms implicitly and concatenate the position and velocity norms for each node $i$ with their respective vectors. Unlike their implementations, we avoid explicit graph construction and do not include edge labels describing atomic bond geometries.
We duplicate all frames $\mathcal{G}^{(t)}\to\{\mathcal{G}^{(t)}\}^{P}$ during dataset initialization, producing a five-fold improvement in throughput compared to previous dataloaders in Table 6.
Table 6: Mean time (seconds) to produce 10 000 batches over 100 benchmark runs. Batch size = 100, 500 samples, $\Delta t=3000$, 500 warmup batches.
Aspirin Ethanol Naphthalene Toluene
EGNO $060{\scriptstyle\pm 0.024} 0.0$$24{\scriptstyle\pm 0.016} 0.0$$56{\scriptstyle\pm 0.024} 0.0$$039{\scriptstyle\pm 0.024}$
ATOM $\textbf{0.005}{\scriptstyle\pm 0.002}$ $\textbf{0.007}{\scriptstyle\pm 0.002}$ $\textbf{0.008}{\scriptstyle\pm 0.004}$ $\textbf{0.006}{\scriptstyle\pm 0.002}$
C.3 Numerical Stability
We evaluate the numerical stability of MD17, RMD17, and TG80. MD17 benzene exhibits substantial center-of-mass drift in Figure 9(a), which is also partially visible in the consistent motion trails shown in Figure 11(a). RMD17 exhibits improved stability, with no center-of-mass drift exceeding $1\times 10^{4}$. TG80 shows the lowest drift of all datasets, and expectedly includes more molecules with high per-step drift (due to more complex sterically hindered geometries).
(a) MD17 molecules are largely consistent, except for benzene, which exhibits substantial drift.
(b) RMD17 molecules are more numerically stable, supporting their use in future benchmarks.
(c) TG80 dataset exhibits the lowest centre-of-mass drift among the evaluated MD datasets.
Figure 9: Comparison of numerical stability across MD17, RMD17, and TG80 datasets. Dashed lines denote the mean centre-of-mass drift and per-step motion; datapoints exceeding two standard deviations are annotated.
C.4 TG80 Generation Algorithm
We first recall the definition of Tanimoto $\operatorname{T}$ similarity between two bit vectors $X,Y$ as
$\operatorname{T}(X,Y)=\frac{|X\cap Y|}{|X\cup Y|},$
which is identical to the definition of the Jaccard similarity in this case (Rogers & Tanimoto, 1960).
To generate TG80, we randomly shuffled the PubChem dataset, then iterated through all compounds until 40 were found that matched the following criteria:
  1. 1.
    Simplified Molecular-input Line-entry System (SMILES) encode a valid molecular structure
  2. 2.
    No more heavy atoms than the corresponding seed molecule
  3. 3.
    Only contain $\{\text{C, H, O, N}\}$ atoms
  4. 4.
    No more than five oxygen atoms
  5. 5.
    No more than three nitrogen atoms
  6. 6.
    No disconnected molecular fragments (e.g., salts)
  7. 7.
    Tanimoto similarity to at least one seed molecule greater than 0.875, less than 0.925
  8. 8.
    Tanimoto similarity to a previously selected molecule is no more than 0.2
This controlled selection procedure generates structurally analogous subsets around each seed molecule whilst preventing convergence to highly similar molecules across different seed groups.
Only 2,488 of the 173 million in the PubChem library satisfied the filtration criteria above. This low yield largely reflects the cumulative effect of criterion 8: as more molecules are added, it becomes harder to find candidates sufficiently dissimilar to all prior selections. Given that the average Tanimoto similarity to our seed set was just 0.1492, the 0.875 threshold was highly selective. Dataset generation code is available at ANONYMIZED.
C.5 Molecular Dynamics Simulations
We present a complete overview of the DFT parameters used to generate MD17 (Chmiela et al., 2017), RMD17 (Christensen & von Lilienfeld, 2020), MD22 (Chmiela et al., 2023), and TG80.
Table 7: An overview of the methodologies used to generate the MD datasets featured.
DFT Functional Dispersion Corrections Basis set Timestep Temperature
MD17 PBE TS NAO 0.5 fs 500K
RMD17 PBE None def2-SVP 0.5 fs 500K
MD22 PBE MBD NAO 1.0 fs 500K
TG80 PBE $\Delta 4$ def2-SVP 1.0 fs 300K
Appendix D Architectural Details
D.1 Fully Equivariant ATOM
To achieve the full equivariance discussed in Figure 7, we employ a canonicalization network approach, which removes Euclidean gauge before learning and then reinstates it afterwards (Kaba et al., 2023). This preserves equivariance of the whole network, even with the use of non-equivariant architectures in the trunk.
We first make data translation equivalent by centering
$\mu=\frac{1}{N}\sum_{i=1}^{N}x_{i},\qquad\bar{x}_{i}=x_{i}-\mu.$ (2)
We then remove rotations by aligning to the second moment
$S=\tfrac{1}{N}\sum_{i=1}^{N}\bar{x}_{i}\bar{x}_{i}^{\top}=\sum_{k=1}^{3}\lambda_{k}e_{k}e_{k}^{\top}\quad(\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}),$ (3)
and choose $e_{1}$ as the principal axis and orthonormalise
$e_{2}\leftarrow\frac{e_{2}-(e_{2}^{\top}e_{1})e_{1}}{\|e_{2}-(e_{2}^{\top}e_{1})e_{1}\|},\qquad e_{3}=e_{1}\times e_{2}.$ (4)
We can then form $Q=[e_{1},e_{2},e_{3}]\in\mathrm{SO}(3)$ and canonicalise
$\tilde{x}_{i}=(x_{i}-\mu)\,Q,\qquad\tilde{v}_{i}=v_{i}\,Q.$ (5)
We fix the eigenvector sign ambiguity using the chirality pseudoscalar $c_{0}=\sum_{i=1}^{N}x_{i}\times v_{i}$ at the reference time (flip $e_{1}$ to satisfy $e_{1}^{\top}c_{0}\geq 0$, then adjust $e_{2},e_{3}$ jointly to keep right-handedness). Let $F$ be an arbitrary trunk acting in the canonical frame; with per-atom canonical outputs $\hat{y}_{i}=F(\{\tilde{x}_{j},\tilde{v}_{j}\}_{j=1}^{N})_{i}$, we decanonicalise by
$y_{i}=\hat{y}_{i}\,Q^{\top}+\mu.$ (6)
This results in exact $\mathrm{SE}(3)$-equivariance (Kaba et al., 2023) and permits non-equivariant trunks.
D.2 Random-Walk Positional Encodings
In the multitask case, we add row-normalized random walk positional encoding (RWPE) to equip ATOM and EGNO with multiscale connectivity features, enhancing their ability to distinguish non-isomorphic graphs (Dwivedi et al., 2022; Ma et al., 2023). We first form a $\varepsilon$-neighborhood graph from our pointclouds as:
$G=(V,E),\quad V=\{i\},\quad E=\{(i,j):\|(x,y,z)_{i}-(x,y,z)_{j}\|_{2}<\varepsilon\}.$ (7)
We set $\varepsilon=1.6$, as covalent bonds typically range from 1.14 Å to 2.0 Å in length (Lobato et al., 2021) and highlight that this construction does not necessitate prior knowledge of the graph structure.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ denote the adjacency matrix of this graph, and let $\mathbf{D}=\text{diag}(\mathbf{A}\mathbf{1})$ represent its degree matrix. We construct the random walk transition matrix as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$ then compute matrix powers of $\mathbf{M}$ up to a maximum walk-length $K$, defining the self-return probabilities for each node as
$p_{i}^{(k)}=\left(\mathbf{M}^{k}\right)_{ii},\quad k=1,\dots,K.$ (8)
These probabilities are collected into vectors $\mathbf{p}_{i}\in\mathbb{R}^{K}$ and concatenated with the phase space to form $\hat{\mathbf{z}}=\left(\mathbf{v}\parallel\mathcal{Z}\parallel\mathbf{p}\right)\in W_{\text{in}}$. Here, the input feature space is redefined as $W_{\text{in}}=V_{\text{in}}\oplus\rho_{0}^{\text{even}}\oplus\left(\rho_{0}^{\text{even}}\otimes\mathbb{R}^{K}\right)$, and the subsequent equivariant maps are modified in kind.
D.3 Value-residual Learning
We employ value-residual learning wherein each transformer block receives the output of the first block via a residual connection to stabilize training and information flow through the network (Zhou et al., 2024b). Inspired by (Jordan, 2024), we add a learned coefficient to weight this residual. Here, $v$ denotes the current block’s value output, and $v_{1}$ represents the initial block’s value. A learnable parameter $\alpha$ is passed through a sigmoid to obtain the weighting coefficient:
$\lambda=\sigma(\alpha).$ (9)
The combined output is then given by:
$v=\lambda\,v+(1-\lambda)\,v_{1}.$ (10)
In practice, we lock the first block’s $\lambda$ value to $0.5$. We report the learned $\lambda$ values in Figure 10.
(a) Aspirin
(b) Uracil
Figure 10: Learned value residuals for MD17 training over 1000 epochs.
D.4 Delta-prediction
When delta-prediction is enabled, as in Figure 7, we incorporate the initial positions $\mathbf{x}$ as a residual term, reformulating the model as an operator that learns a displacement field rather than predicting absolute positions. We express this as:
$\mathbf{x}^{\dagger}=\operatorname{Project}(\mathbf{x}_{\text{out}})+\mathbf{x}.$ (11)
Although this approach is implemented in both EGNN and EGNO, we found it was disabled by default in the codebase of the latter (Garcia Satorras et al., 2021; Xu et al., 2024). Based on empirical results from our ablations, Figure 7, we argue there is sufficient evidence to discourage the use of delta-prediction in neural operator-based molecular dynamics simulations.
Appendix E Further Experiments
E.1 Results on tail discretization
We find the performance of both EGNO and ATOM on MD17 with tail discretization remains similar to the performance under uniform discretization discussed in table 1.
Table 8: EGNO and ATOM with final frame sampling
Aspirin Benzene Ethanol Malonaldehyde Naphthalene Salicylic Toluene Uracil
EGNO $9.66{\scriptstyle\pm 0.12}$ $39.09{\scriptstyle\pm 2.35}$ $4.57{\scriptstyle\pm 0.01}$ $\mathbf{12.92}{\scriptstyle\pm 0.00}$ $0.39{\scriptstyle\pm 0.00}$ $0.88{\scriptstyle\pm 0.01}$ $10.99{\scriptstyle\pm 0.00}$ $0.60{\scriptstyle\pm 0.00}$
ATOM $\mathbf{6.38}{\scriptstyle\pm 0.17}$ $\mathbf{39.03}{\scriptstyle\pm 3.32}$ $\mathbf{3.62}{\scriptstyle\pm 0.08}$ $15.26{\scriptstyle\pm 0.65}$ $\mathbf{0.39}{\scriptstyle\pm 0.00}$ $\mathbf{0.83}{\scriptstyle\pm 0.01}$ $\mathbf{5.26}{\scriptstyle\pm 0.79}$ $\mathbf{0.55}{\scriptstyle\pm 0.00}$
Gap $33.97\% +$$0.15\% +$$20.85\%$ $18.06\%$ $1.62\% +$$4.75\% +$$52.13\% +$$9.28\%$
EGNO $9.66{\scriptstyle\pm 0.11}$ $39.15{\scriptstyle\pm 2.28}$ $4.57{\scriptstyle\pm 0.01}$ $\mathbf{12.92}{\scriptstyle\pm 0.01}$ $0.39{\scriptstyle\pm 0.00}$ $0.88{\scriptstyle\pm 0.01}$ $10.99{\scriptstyle\pm 0.00}$ $0.60{\scriptstyle\pm 0.00}$
ATOM $\mathbf{6.38}{\scriptstyle\pm 0.17}$ $\mathbf{39.03}{\scriptstyle\pm 3.35}$ $\mathbf{3.63}{\scriptstyle\pm 0.08}$ $15.21{\scriptstyle\pm 0.60}$ $\mathbf{0.38}{\scriptstyle\pm 0.00}$ $\mathbf{0.83}{\scriptstyle\pm 0.01}$ $\mathbf{5.27}{\scriptstyle\pm 0.79}$ $\mathbf{0.55}{\scriptstyle\pm 0.00}$
Gap $33.91\% +$$0.30\% +$$20.66\%$ $17.71\%$ $1.82\% +$$5.02\% +$$52.08\% +$$9.44\%$
E.2 Revised MD17 Dataset
We reach performance parity with EGNO on RMD17, shown in Table 9.
Table 9: EGNO and ATOM with final frame sampling
Azobenzene Ethanol Malonaldehyde Naphthalene Paracetamol Salicylic Toluene Uracil
EGNO $8.96{\scriptstyle\pm 0.03}$ $\mathbf{23.26}{\scriptstyle\pm 0.01}$ $\mathbf{40.11}{\scriptstyle\pm 0.05}$ $1.42{\scriptstyle\pm 0.00}$ $\mathbf{28.08}{\scriptstyle\pm 0.01}$ $1.06{\scriptstyle\pm 0.01}$ $\mathbf{28.28}{\scriptstyle\pm 0.01}$ $0.88{\scriptstyle\pm 0.00}$
ATOM $\mathbf{8.88}{\scriptstyle\pm 0.05}$ $23.49{\scriptstyle\pm 0.14}$ $40.29{\scriptstyle\pm 0.13}$ $\mathbf{1.36}{\scriptstyle\pm 0.00}$ $30.12{\scriptstyle\pm 0.87}$ $\mathbf{1.03}{\scriptstyle\pm 0.00}$ $28.56{\scriptstyle\pm 0.04}$ $\mathbf{0.86}{\scriptstyle\pm 0.00}$
Gap $0\% 0.99-0$$9\%$ $0.45\%$ $3.93\%$ $7.26\%$ $3.10\%$ $0.99\%$ $1.90\%$
EGNO $8.51{\scriptstyle\pm 0.03}$ $\mathbf{23.61}{\scriptstyle\pm 0.03}$ $\mathbf{40.32}{\scriptstyle\pm 0.08}$ $1.42{\scriptstyle\pm 0.00}$ $\mathbf{28.01}{\scriptstyle\pm 0.02}$ $1.07{\scriptstyle\pm 0.01}$ $\mathbf{28.23}{\scriptstyle\pm 0.00}$ $0.87{\scriptstyle\pm 0.00}$
ATOM $\mathbf{8.38}{\scriptstyle\pm 0.05}$ $23.90{\scriptstyle\pm 0.15}$ $40.67{\scriptstyle\pm 0.17}$ $\mathbf{1.36}{\scriptstyle\pm 0.00}$ $30.03{\scriptstyle\pm 0.78}$ $\mathbf{1.04}{\scriptstyle\pm 0.00}$ $28.58{\scriptstyle\pm 0.05}$ $\mathbf{0.85}{\scriptstyle\pm 0.00}$
Gap $1.47\%$ $1.27\%$ $0.88\%$ $4.39\%$ $7.21\%$ $2.78\%$ $1.23\%$ $2.00\%$
E.3 Random-split cross-validation on TG80.
For completeness, we report multitask results under compound-level random cross-validation, where compounds are randomly assigned to the train, validation, and test sets. Relative to the more challenging out-of-domain (UMAP-based) split in Table 3, EGNO is comparatively stronger; nevertheless, ATOM maintains a consistent lead across folds, with mean improvements of 24.43% on S2S and 23.93% on S2T.
Table 10: S2S MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
Cluster 1 Cluster 2 Cluster 3 Cluster 4 Cluster 5
OOD EGNO $71.83{\scriptstyle\pm 0.00}$ $76.92{\scriptstyle\pm 0.00}$ $68.99{\scriptstyle\pm 0.00}$ $101.27{\scriptstyle\pm 0.00}$ $83.20{\scriptstyle\pm 0.00}$
ATOM $\mathbf{53.93}{\scriptstyle\pm 0.00}$ $\mathbf{62.40}{\scriptstyle\pm 0.00}$ $\mathbf{49.37}{\scriptstyle\pm 0.00}$ $\mathbf{70.75}{\scriptstyle\pm 0.00}$ $\mathbf{66.75}{\scriptstyle\pm 0.00}$
Gap $24.92\% +$$18.88\% +$$28.45\% +$$30.14\% +$$19.77\%$
Table 11: S2T MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
Cluster 1 Cluster 2 Cluster 3 Cluster 4 Cluster 5
OOD EGNO $63.23{\scriptstyle\pm 0.00}$ $64.49{\scriptstyle\pm 0.00}$ $59.18{\scriptstyle\pm 0.00}$ $85.87{\scriptstyle\pm 0.00}$ $69.46{\scriptstyle\pm 0.00}$
ATOM $\mathbf{46.09}{\scriptstyle\pm 0.00}$ $\mathbf{54.47}{\scriptstyle\pm 0.00}$ $\mathbf{42.90}{\scriptstyle\pm 0.00}$ $\mathbf{55.64}{\scriptstyle\pm 0.00}$ $\mathbf{59.55}{\scriptstyle\pm 0.00}$
Gap $27.10\% +$$15.54\% +$$27.51\% +$$35.21\% +$$14.28\%$
E.4 Multitask S2S results on TG80 under UMAP cluster cross-validation.
Table 12: S2S MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
Cluster 1 Cluster 2 Cluster 3 Cluster 4 Cluster 5
ID EGNO $51.98{\scriptstyle\pm 0.81}$ $95.86{\scriptstyle\pm 0.53}$ $142.51{\scriptstyle\pm 0.58}$ $155.25{\scriptstyle\pm 0.67}$ $109.25{\scriptstyle\pm 0.24}$
ATOM $\mathbf{15.49}{\scriptstyle\pm 1.04}$ $\mathbf{26.55}{\scriptstyle\pm 2.13}$ $\mathbf{28.74}{\scriptstyle\pm 2.40}$ $\mathbf{29.81}{\scriptstyle\pm 2.72}$ $\mathbf{26.33}{\scriptstyle\pm 1.98}$
Gap (%) $70.20\%$ $72.30\%$ $79.83\%$ $80.80\%$ $75.90\%$
OOD EGNO $52.90{\scriptstyle\pm 0.72}$ $114.14{\scriptstyle\pm 13.21}$ $149.99{\scriptstyle\pm 0.34}$ $63.47{\scriptstyle\pm 1.00} 11$$12.36{\scriptstyle\pm 1.90}$
EGNN-S $52.39{\scriptstyle\pm 0.40}$ $16512.07{\scriptstyle\pm 12314.09}$ $149.41{\scriptstyle\pm 0.94}$ $663.54{\scriptstyle\pm 865.23}$ $111.08{\scriptstyle\pm 0.62}$
EGNN-R $52.08{\scriptstyle\pm 0.79}$ $\mathbf{108.89}{\scriptstyle\pm 1.60}$ $148.67{\scriptstyle\pm 0.73}$ $163.27{\scriptstyle\pm 0.16}$ $109.94{\scriptstyle\pm 0.31}$
ATOM $\mathbf{41.97}{\scriptstyle\pm 1.24}$ $127.95{\scriptstyle\pm 122.67}$ $\mathbf{74.53}{\scriptstyle\pm 4.82}$ $\mathbf{80.95}{\scriptstyle\pm 1.21}$ $\mathbf{58.26}{\scriptstyle\pm 1.68}$
Gap (%) $19.41\%$ $17.50\%$ $49.87\%$ $50.42\%$ $47.01\%$
E.5 Explicit Hydrogen Representation
We hypothesized that our model’s underperformance relative to EGNO in predicting malonaldehyde was due to the omission of explicit hydrogens, which limits its ability to capture electron delocalization effects involving hydroxyl hydrogens. We tested two explicit hydrogen methods:
  1. 1.
    Including all hydrogens with gradients computed for all atoms during training
  2. 2.
    Including all hydrogens but computing gradients only for heavy atoms
Contrary to conventional MD practice, neither method improved heavy-atom test loss.
Table 13: ATOM S2T MSE with implicit hydrogens (ATOM baseline) and two explicit hydrogen approaches.
Aspirin Malonaldehyde
Implicit hydrogens $6.52{\scriptstyle\pm 0.08}$ $13.51{\scriptstyle\pm 0.10}$
Explicit hydrogens 1 $7.48{\scriptstyle\pm 0.11}$ $15.15{\scriptstyle\pm 0.22}$
Explicit hydrogens 2 $6.96{\scriptstyle\pm 0.20}$ $13.52{\scriptstyle\pm 0.10}$
Appendix F Experimental Details
F.1 Software and Hardware Details
All experiments were conducted using Python 3.12, NumPy 2.2.1 (Harris et al., 2020), PyTorch 2.5.1 (Paszke et al., 2019), e3nn 0.5.6 (Geiger & Smidt, 2022) and PyTorch Optimizer 3.5.0 (Kim, 2021). We use RDKit 2024.9.6 (Landrum et al., 2025) and PubChemPy 1.0.4 in the construction of TG80. All single-task training was performed on an NVIDIA RTX 5080 (16 GB) with CUDA 12.4, running on Ubuntu 24.04.
F.2 Computational Costs of Model Training
We roughly wall-clock normalised our ATOM and EGNO parameter counts, resulting in respective learnable parameter counts of 754 468 and 335 770. Despite this, ATOM still trained approximately 20% faster in single-task learning on MD17 and TG80 over three runs.
Table 14: Compute cost of single-task training on all MD17 molecules over 1000 epochs. Both ATOM (335 770 params) and EGNO (754 468 params) are under torch.compile on a Titan V.
Model Metric Azobenzene Ethanol Malonaldehyde Naphthalene Paracetamol Salicylic Toluene Uracil
EGNO Time (mins) $4.09\pm 0.15$ $3.62\pm 0.42$ $3.65\pm 0.01$ $5.01\pm 0.02$ $9.02\pm 1.22$ $5.16\pm 0.03$ $3.93\pm 0.02$ $4.35\pm 0.04$
Total FLOPS ($\times 10^{12}$) $3681.24$ $3257.70$ $3282.09$ $4513.37$ $8114.44$ $4641.50$ $3539.92$ $3915.98$
Epochs/min $244.48$ $276.27$ $274.22$ $199.41$ $110.91$ $193.90$ $254.24$ $229.83$
ATOM Time (mins) $5.81\pm 0.02$ $5.79\pm 0.06$ $5.79\pm 0.00$ $5.86\pm 0.01$ $5.89\pm 0.02$ $5.85\pm 0.01$ $5.81\pm 0.01$ $5.83\pm 0.02$
Total FLOPS ($\times 10^{12}$) $5226.49$ $5212.53$ $5213.50$ $5271.19$ $5297.84$ $5263.76$ $5224.91$ $5247.33$
Epochs/min $172.20$ $172.66$ $172.63$ $170.74$ $169.88$ $170.98$ $172.25$ $171.52$
Total FLOPS Reduction (%) $41.98\%$ $60.01\%$ $58.85\%$ $16.79\%$ $34.71\%$ $13.41\%$ $47.60\%$ $34.00\%$
In multitask training on TG80, our upsized ATOM model contained 3 557 840 parameters, compared to XXX for EGNO. Despite this, ATOM only trained between 5% and 30% slower than EGNO. This is perhaps unsurprising given the much higher FLOPS-utilization of the transformer architecture upon which ATOM is based.
Table 15: Compute cost of single task training on five TG80 molecules over 1000 epochs. Both ATOM (335 770 params) and EGNO (754 468 params) are under torch.compile on a Titan V.
Model Fold1 Fold2 Fold3 Fold4 Fold5
EGNO Time (mins) $9.61{\scriptstyle\pm 1.21}$ $8.56{\scriptstyle\pm 0.06}$ $9.04{\scriptstyle\pm 0.11}$ $9.31{\scriptstyle\pm 0.20}$ $8.98{\scriptstyle\pm 0.01}$
Total FLOPS ($\times 10^{12}$) 8645.66 7703.33 8136.98 8378.06 8084.98
Epochs/min $104.10$ $116.83$ $110.61$ $107.42$ $111.32$
ATOM Time (mins) $10.16{\scriptstyle\pm 0.49}$ $0.55{\scriptstyle\pm 0.02} 11$$1.62{\scriptstyle\pm 0.41}$ $12.38{\scriptstyle\pm 0.12}$ $10.39{\scriptstyle\pm 0.41}$
Total FLOPS ($\times 10^{12}$) 9140.57 9497.79 10455.34 11141.73 9355.37
Epochs/min $98.46$ $94.76$ $86.08$ $80.78$ $96.20$
Total FLOPS Reduction (%) $5.72\%$ $23.29\%$ $28.49\%$ $32.99\%$ $15.71\%$
F.3 ATOM Hyperparameters
We employ the same dataset splitting and discretization parameters reported in Xu et al. (2024) for the MD17. We set the batch size to $192$, use the AdamW-AMSGrad optimizer (Loshchilov & Hutter, 2017) with an $\epsilon$ of $1\times 10^{-10}$ to avoid instability associated with the small gradients produced by zero-initialised weight matrices in early training (Jordan et al., 2025). During multitask training, we reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). We present a complete overview of our hyperparameters in Table 16.
Table 16: Hyperparameters for ATOM. MD17 hyperparameters are shared across all molecules unless otherwise noted.
Module MD17, RMD17, TG80 TG80 Multitask
Training
Batch size $192$ $192$
Epochs $1000$ $250$
Max grad norm $1.0$ $1.0$
Label noise $\sigma$ $0.1$ $0.1$
$\Delta t$ $3000$ 10 00010\ 000
Timesteps $P$ 88 88
Train/Val/Test (500,3 000,3 000)(500,3\,000,3\,000) (6 500,13 000,13 000)(6\,500,13\,000,13\,000)
RWPE length 8 8
Optimiser
Optimiser type AdamW-AMSGrad Muon
Learning rate $\times 10^{-3} 1×1$$\times 10^{-3}$
$\beta_{1},\beta_{2}$ $0.9,0.999$ $(0.9,0.999)$
Weight decay $\times 10^{-5} 1×1$$\times 10^{-5}$
$\epsilon$ $\times 10^{-10} 1×1$$\times 10^{-5}$
Model
Embedding dim $128$ $256$
No. layers 55 66
No. attention heads 88 88
No. output heads 11 88
Attention dropout $0.2$ $0.2$
RoPE frequency $1000$ $1000$
MLP layers 22 22
MLP activation SwiGLU SwiGLU
MLP dropout $0.0$ $0.0$
Norm type RMS norm RMS norm
Learnable value residuals True True
F.4 EGNO Hyperparameters and Experimental Details
We generated the EGNO results reported in Table 1 with the same discretization parameters and hyperparameters as used in their experiments. We reduce the number of epochs from 10 00010\ 000 to 2 5002\ 500, use a batch size of 192 with the AdamW-AMSGrad optimizer (Loshchilov & Hutter, 2017), and select the best validation loss epoch for testing. In the multitask case, we further reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). Complete hyperparameters are displayed in Table 17.
Table 17: Hyperparameter values for EGNO across each benchmark dataset.
Module MD17, RMD17, TG80 TG80 Multitask
Training
Batch size $192$ $192$
Epochs $2500$ $250$
Max grad norm Uncapped Uncapped
Label noise $\sigma$ $0.1$ $0.1$
$\Delta t$ $3000$ 10 00010\ 000
Timesteps $P$ 88 88
Train/Val/Test (500,3 000,3 000)(500,3\,000,3\,000) (6 500,13 000,13 000)(6\,500,13\,000,13\,000)
RWPE length 8 8
Optimiser
Optimiser type AdamW-AMSGrad Muon
Learning rate $\times 10^{-3} 1×1$$\times 10^{-3}$
$\beta_{1},\beta_{2}$ $(0.9,0.999)$ $(0.9,0.999)$
Weight decay $\times 10^{-5} 1×1$$\times 10^{-5}$
$\epsilon$ $\times 10^{-10} 1×1$$\times 10^{-5}$
Scheduler
Scheduler type StepLR StepLR
Step size 2500 2500
$\gamma$ 0.5 0.5
Model
Embedding dim $64$ $64$
No. EGNO layers 55 55
Temporal convolution activation LeakyRELU LeakyRELU
MLP layers 22 22
MLP activation SiLU SiLU
MLP dropout 0 0
Time embedding dim 32 32
Fourier modes 2 2
Appendix G Propositions and Proofs
G.1 Kernel Integral form of Cross-attention
Proposition G.1.
The cross-attention is equivalent to a kernel integral operator, i.e., $\operatorname{softmax}(\operatorname{T-RoPE}(\mathbf{Q})\operatorname{T-RoPE}(\mathbf{K}_{i})^{\top}/\sqrt{d_{h}})\mathbf{V}_{i}=\int\kappa_{i}(\mathbf{z},\mathbf{x})v_{i}(x)d\mu_{N}(\mathbf{x})$, where $\kappa_{i}$ denotes the kernel induced by softmax function, $v_{i}(\mathbf{x})$ denotes the values as a function of $\mathbf{x}$, and $\mu_{N}$ denotes the empirical measure supported on $\{\mathbf{x}_{j}\}_{j=1}^{N}$.
Proof of Proposition G.1.
Following (Gao et al., 2024) we may view our attention as a kernel integral transform by considering $\mathbf{x}_{i}$ as being sampled from the continuum domain $\Omega\subset\mathbb{R}^{3}$ for which we define the empirical measure with support on $\{\mathbf{x}_{j}\}_{j=1}^{N}\subset\Omega$:
${\mu_{N}}(\mathbf{x})=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{x}_{j}},\qquad\int_{\Omega}g(\mathbf{x})d{\mu_{N}}(\mathbf{x})=\frac{1}{N}\sum_{j=1}^{N}g(\mathbf{x}_{j})$ (12)
where $\delta$ is the Dirac delta function “selecting” the values at $\mathbf{x}_{j}$. Given T-RoPE-rotated query and key maps $\tilde{q}_{\theta}(\mathbf{z})=R_{p(\mathbf{z})}q_{\theta}(\mathbf{z}_{j})$, $\tilde{k}_{i}(\mathbf{x}_{j})=R_{p(\mathbf{x}_{j})}k_{\theta,i}(\mathbf{x}_{j})$ we form the data-dependent kernel for feature $F$:
$\kappa_{\theta,i}(\mathbf{z},\mathbf{x}_{j})=\frac{\exp\bigl(\langle\tilde{q}(\mathbf{z}),\,\tilde{k}_{i}(\mathbf{x}_{j})\rangle\;/\;\sqrt{d_{h}}\bigr)}{\displaystyle\int_{\Omega}\exp\bigl(\langle\tilde{q}(\mathbf{z}_{j}),\,\tilde{k}_{i}(\mathbf{x}^{\prime})\rangle\;/\;\sqrt{d_{h}}\bigr)\,d\mu_{N}(\mathbf{x}^{\prime})}\,.$ (13)
Thus, for any $F\in\mathcal{F}$ we may represent our cross-attention as the kernel integral operator:
$\bigl(\mathcal{K}_{\theta,i}\mathbf{v}_{j}\bigr)(\mathbf{z})=\int_{\Omega}\kappa_{\theta,i}\bigl(\mathbf{z},\mathbf{x}\bigr)\,\mathbf{v}_{i}(\mathbf{x})\,\mathrm{d}\mu_{N}(\mathbf{x}),\qquad\int_{\Omega}\kappa_{\theta,i}(\mathbf{z},\mathbf{x})\,d\mu_{N}(\mathbf{x})=1,$ (14)
which is row-stochastic under the measure in Equation 12. ∎
We remark that the kernel fails to satisfy global Lipschitz continuity (Delattre et al., 2023), unlike FNO (Li et al., 2021), and certain generalization theorems fail as a result (Le & Dik, 2024).
Appendix H Trajectory Samples
(a) MD17 trajectories
(b) RMD17 trajectories
Figure 11: 3000 steps MD trajectories from the MD17 and RMD17 datasets.
Figure 12: 3000-step MD trajectories from TG80. Molecules generated by our dataset expansion algorithm are named according to their seed molecule and the order of their selection.